# LLMInit: A Free Lunch from Large Language Models for Selective Initialization of Recommendation

### v1

@@
Abstract.
Collaborative filtering models, particularly graph-based approaches, have demonstrated strong performance in capturing user-item interactions for recommendation systems. However, they continue to struggle in cold-start and data-sparse scenarios. The emergence of large language models (LLMs) like GPT and LLaMA presents new possibilities for enhancing recommendation performance, especially in cold-start settings. Despite their promise, LLMs pose challenges related to scalability and efficiency due to their high computational demands and limited ability to model complex user-item relationships effectively. In this work, we introduce a novel perspective on leveraging LLMs for CF model initialization. Through experiments, we uncover an embedding collapse issue when scaling CF models to larger embedding dimensions. To effectively harness large-scale LLM embeddings, we propose innovative selective initialization strategies utilizing random, uniform, and variance-based index sampling. Our comprehensive evaluation on multiple real-world datasets demonstrates significant performance gains across various CF models while maintaining a lower computational cost compared to existing LLM-based recommendation approaches.
Collaborative Filtering, Recommendation, Large Language Models
doi: XXXXXXX.XXXXXXXccs: Information systems Recommender systemsccs: Information systems Collaborative filtering
1. Introduction
Recommender systems (RecSys), particularly collaborative filtering (CF), play a pivotal role in many online platforms and web applications  (Koren et al., 2021; Fan et al., 2019; Zou et al., 2024), aiding users in navigating vast information by offering personalized suggestions. Among the various techniques, collaborative filtering (Wang et al., 2019; He et al., 2020; Xu et al., 2023; Zhang et al., 2024a; Xu et al., 2025) stands out for its effectiveness in predicting user preferences by analyzing past user-item interactions. A key principle among these models is learning randomly initialized user/item representations with carefully designed loss functions to capture each user’s preference (Rendle et al., 2012). However, such CF approaches heavily rely on the observed user-item interactions to optimize the CF embeddings from random initialization and perform poorly in real-world applications where interaction data is sparse and users/items are cold-start.
More recently, the rapid advancement of large language models (LLMs) such as GPT-4  (Achiam et al., 2023) and LLaMa  (Touvron et al., 2023), which have exhibited remarkable proficiency in understanding and processing textual information, have piqued the interest of researchers seeking to explore their potential in improving recommender systems beyond traditional approaches  (Geng et al., 2022; Yuan et al., 2023; Li et al., 2023). One of the most promising directions in this exploration involves adapting LLMs as recommender systems through prompt engineering and tailored instruction tuning. Such integration of LLMs enables the RecSys to better understand complex user preferences, and interpret contextual information, thereby surpassing the traditional models in cold-start settings.
@@ 1. Introduction
Figure 1. An investigation of the embedding collapse issue of CF model LightGCN (He et al., 2020) in two Amazon datasets. Increasing embedding size with exponentially growing parameters finally leads to performance degradation.
@@
Despite the inspiring progress made in LLMs for recommendations, their adoption in online deployments faces significant challenges related to efficiency and scalability. These issues stem from the inherently time-consuming and computationally intensive nature of LLMs. Such problems can be further exaggerated in real-world applications of large-scale users and items due to the enlarged vocabulary tokens and in-context input. Moreover, LLMs fall short of complex CF interaction understanding owing to the next-token prediction pipeline and the constraints of a maximum token limit (e.g., 4096 tokens in LLaMA) to incorporate the whole CF interactions. This limitation hampers their ability to capture and effectively model global user-item dependencies as in CF models, leading to poor performance in the full-ranking (Hou et al., 2024) and warm-start settings (Bao et al., 2023). Therefore, we raise a critical research question:
How can we harness the power of LLMs for recommendation systems in an effective and scalable manner?
@@
To answer the above question, we propose to utilize LLMs for CF-based RecSys from a new perspective by taking pretrained large language models as a free lunch for the embedding initialization of CF models. One trivial way is to directly utilize the LLMs to generate embeddings with substantial textual information. However, storing the large-scale embeddings from LLMs for each user and item is not efficient nor scalable, especially when the platform scales to millions of users and items. Furthermore, as illustrated in Figure 1, unlike the scaling laws observed in NLP models (Kaplan et al., 2020), the performance of the recommendation model LightGCN (He et al., 2020) deteriorates even when the numbers of parameters scale up exponentially. This behavior resembles the embedding collapse issue (Guo et al., [n. d.]) observed in click-through rate prediction models (Guo et al., 2017; Wang et al., 2021). These findings suggest that significantly smaller embedding sizes (e.g., 128) are more suitable for recommendation tasks. In contrast, the top-10 models in the LLM embedding benchmark MTEB  [^1] https://huggingface.co/spaces/mteb/leaderboard (Muennighoff et al., 2022), pretrained for various NLP tasks, use an average large embedding size of 4,506.
In this paper, we propose LLMInit, a novel initialization framework to bridge the embedding gaps between LLMs and CF models for effective and scalable recommendations. Specifically, we propose three strategies for selecting and transferring essential knowledge from LLM embeddings to CF models. For item embedding initialization, we explore the random selection, uniform selection, and variance-based selection to sample embedding index dimensions from LLM-generated representations. On the user side, where contextual information is frequently missing, we aggregate embeddings via a pooling operation over item neighborhoods. These approaches enable small CF models to selectively inherit the rich semantics of large language model embeddings while maintaining efficiency and scalability during the training and inference stage. Extensive experiments demonstrate the efficiency of LLMInit when applied to several state-of-the-art (SOTA) CF models and compared to other LLM-based models. We summarize our contributions as follows:
@@
  • Methodologically, we devise three types of selective initialization strategies to inherit the rich world knowledge from pretrained LLMs into lightweight CF models for scalable recommendations.
  • Empirically, we implement LLMInit in a plug-and-play manner on various SOTA CF models, achieving significant performance and efficiency gains, especially compared to LLM-based RecSys.
2. Preliminaries
2.1. Problem Definition
@@
Figure 2. An Illustration of LLMInit framework including contextual LLM input, three types of selective item embedding initialization, and the user embedding aggregation.
2.2. Related Work
LLMs for Recommendation. Recent advancements in this domain can be categorized into three paradigms (Wu et al., 2024): extracting embeddings for sequential recommendation (Qiu et al., 2021; Hou et al., 2022), generating semantic tokens to capture user preferences (Li et al., 2023; Xi et al., 2024), and directly adapting LLMs as RecSys through prompting (Hou et al., 2024; Liang et al., 2025) or tuning (Geng et al., 2022). However, they all fail to capture the intricate CF relationship in the user-item bipartite graphs as they either only focus on sequential order patterns (Qiu et al., 2021; Hou et al., 2022) or semantic understandings (Li et al., 2023; Xi et al., 2024; Hou et al., 2024; Zhang et al., 2023; Geng et al., 2022). In addition, they face notable challenges in efficiency and scalability due to the computational demands of LLMs, large-scale embeddings, overloaded vocabulary tokens, and limited in-context inputs in real-world scenarios. In contrast, our work LLMInit, proposes a novel paradigm to leverage LLMs in a scalable manner and addresses ineffective interaction modeling by resorting to CF models.
Table 1. Performance comparison of different CF models on four benchmark datasets in terms of Recall@10 and NDCG@10.
Method Beauty Toys-Games Tools-Home Office-Products
Recall@10 NDCG@10 Recall@10 NDCG@10 Recall@10 NDCG@10 Recall@10 NDCG@10
LightGCN (He et al., 2020) 0.0910 0.0432 0.0775 0.0360 0.0574 0.0283 0.0745 0.0365
+LLMInit-Rand 0.0960 (+5.5%) 0.0467 (+8.1%) 0.0805 (+3.9%) 0.0387 (+7.50%) 0.0612 (+6.6%) 0.0313 (+10.6%) 0.0773 (+3.8%) 0.0387 (+6.0%)
+LLMInit-Uni 0.1006 (+10.6%) 0.0469 (+8.6%) 0.0806 (+4.0%) 0.0388 (+7.8%) 0.0633 (+10.3%) 0.0319 (+12.7%) 0.0791 (+6.2%) 0.0395 (+8.2%)
+LLMInit-Var 0.1019 (+12.0%) 0.0485 (+12.3%) 0.0808 (+4.3%) 0.0389 (+8.1%) 0.0633 (+10.3%) 0.0317 (+12.0%) 0.0816 (+9.5%) 0.0414 (+13.4%)
SGL (Wu et al., 2021) 0.1017 0.0474 0.0832 0.0380 0.0580 0.0284 0.0669 0.0297
+LLMInit-Rand 0.1069 (+5.1%) 0.0520 (+9.7%) 0.0885 (+6.4%) 0.0418 (+10.0%) 0.0692 (+19.3%) 0.0337 (+18.7%) 0.0810 (+21.1%) 0.0426 (+43.4%)
+LLMInit-Uni 0.1101 (+8.3%) 0.0513 (+8.2%) 0.0920 (+10.6%) 0.0424 (+11.6%) 0.0676 (+16.6%) 0.0333 (+17.3%) 0.0773 (+15.6%) 0.0350 (+17.9%)
+LLMInit-Var 0.1106 (+8.8%) 0.0530 (+11.8%) 0.0927 (+11.4%) 0.0427 (+12.4%) 0.0686 (+18.3%) 0.0339 (+19.4%) 0.0794 (+18.7%) 0.0421 (+41.8%)
SGCL (Zhang et al., 2024b) 0.1027 0.0499 0.0828 0.0382 0.0585 0.0294 0.0647 0.0298
+LLMInit-Rand 0.1094 (+6.5%) 0.0512 (+2.6%) 0.0929 (+12.2%) 0.0418 (+9.4%) 0.0651 (+11.3%) 0.0326 (+10.9%) 0.0770 (+19.0%) 0.0365 (+22.5%)
+LLMInit-Uni 0.1115 (+8.6%) 0.0513 (+2.8%) 0.0923 (+11.5%) 0.0422 (+10.5%) 0.0650 (+11.1%) 0.0327 (+11.2%) 0.0742 (+14.7%) 0.0354 (+18.8%)
+LLMInit-Var 0.1104 (+7.5%) 0.0522 (+4.6%) 0.0941 (+13.6%) 0.0421 (+10.2%) 0.0646 (+10.4%) 0.0327 (+11.2%) 0.0776 (+19.9%) 0.0366 (+22.8%)
3. Method
@@
3.1. Selective Item Embedding Initialization
The CF item representation space can be considered a homomorphic projection of the sophisticated world-knowledge representation space in large language models (Sheng et al., 2024). Motivated by this, we propose to distill and selectively utilize the embeddings from LLMs to initialize the embeddings for CF recommendation models.
Formally, $\mathbf{v}\in\mathbb{R}^{N}$ represents the item embedding vector generated by the pre-trained LLMs. Let $K$ denotes the desired embedding dimensionality after selection. Then $\mathcal{I}\subseteq\{0,1,\dots,N-1\}$ is the selected indices. The resulting $K$-dimensional embedding is: $\mathbf{v}_{\text{selected}}=\{\mathbf{v}[i]:i\in\mathcal{I}\}.$ Here, we propose three novel initialization strategies to mitigate the inefficiency and embedding collapse issue for directly adopting the LLMs embeddings as follows.
3.1.1. Uniform Selection (LLMInit-Uni)
@@
=$\mathcal{I}=\{k\cdot\lfloor N/K\rfloor:k\in\{0,1,\dots,K-1\}\}.$
3.1.2. Random Selection (LLMInit-Rand)
@@
3.1.3. Variance Selection (LLMInit-Var)
@@
3.2. Aggregated User Embedding Initialization
@@
(1) $\mathbf{u}_{i}=\sum_{j\in N_{i}}\frac{1}{{|N_{i}|}}\mathbf{v}_{j},$
@@
4. Experiments
4.1. Experimental Set Up
Table 2. The statistics of the datasets.
Dataset # Users # Items # Interactions Sparsity
Beauty 10,554 6,087 94,148 99.8534%
Toys-Games 11,269 7,310 95,420 99.8842%
Tools-Home 9,246 6,199 73,250 99.8722%
Office-Products 3,261 1,584 36,544 99.2925%
We follow the 5-core setting in (Zhang et al., 2024b) and experiment on four Amazon  [^2] https://jmcauley.ucsd.edu/data/amazon/links.html real-world datasets: Beauty, Toys-Games, Tools-Home, and Office-Products, as summarized in Table  2. We use a leave-one-out strategy for data splitting and NDCG@10 and Recall@10 metrics for evaluation. For the CF baselines, we select LightGCN (He et al., 2020) as a representative of graph models, SGL (Wu et al., 2021) for self-supervised contrastive learning approaches, and SGCL (Zhang et al., 2024b) for the supervised contrastive learning paradigm. For LLM-based baselines, we use MoRec (Yuan et al., 2023) , LLMRank (Hou et al., 2024), and LLMRec (Wei et al., 2024). For fair comparisons, we set embedding size $K=128$ for all methods and LLM as MPNet (Song et al., 2020) for embedding generation unless specified.
4.2. Performance Comparison
The experimental results demonstrate that the proposed LLMInit strategies significantly enhance performance across all datasets compared to traditional baselines. Among all LLMInit strategies, LLMInit-Var consistently achieves the best results, leveraging variance-based initialization to capture CF and semantic nuances effectively. While LLMInit-Rand and LLMInit-Uni also outperform the baselines, the variance-based strategy emerges as the most robust and adaptable approach. Notably, in the Office-Products dataset, LLMInit delivers an average over 20% boost in NDCG@10 over SGL and SGCL. These findings underscore the robustness and effectiveness of LLMInit strategies, particularly LLMInit-Var, in enhancing recommendation and handling diverse datasets.
4.3. Efficiency Comparison of LLM-based RecSys
Table 3. Comparisons of computation cost and performance with LLM-based RecSys in full-ranking settings.
@@ 4.3. Efficiency Comparison of LLM-based RecSys
MoRec (Yuan et al., 2023) 0.0857 0.0397 13M 0.0793 0.0362 14M
LLMRank (Hou et al., 2024) 7B 7B
LLMRec (Wei et al., 2024) 0.0974 0.0472 7B+13M 0.082 0.0389 7B+14M
LLMInit 0.1104 0.0522 0.2M 0.0941 0.0421 0.2M

### v2

@@
Abstract
Collaborative filtering (CF) is widely adopted in industrial recommender systems (RecSys) for modeling user-item interactions across numerous applications, but often struggles with cold-start and data-sparse scenarios. Recent advancements in pre-trained large language models (LLMs) with rich semantic knowledge, offer promising solutions to these challenges. However, deploying LLMs at scale is hindered by their significant computational demands and latency. In this paper, we propose a novel and scalable LLM-RecSys framework, LLMInit, designed to integrate pretrained LLM embeddings into CF models through selective initialization strategies. Specifically, we identify the embedding collapse issue observed when CF models scale and match the large embedding sizes in LLMs and avoid the problem by introducing efficient sampling methods, including, random, uniform, and variance-based selections. Comprehensive experiments conducted on multiple real-world datasets demonstrate that LLMInit significantly improves recommendation performance while maintaining low computational costs, offering a practical and scalable solution for industrial applications. To facilitate industry adoption and promote future research, we provide open-source access to our implementation at https://github.com/DavidZWZ/LLMInit.
LLMInit: A Free Lunch from Large Language Models
for Selective Initialization of Recommendation
Weizhi Zhang1, Liangwei Yang2thanks: Corresponding author: liangwei.yang@salesforce.com, Wooseong Yang1, Henry Peng Zou1, Yuqing Liu1, Ke Xu1, Sourav Medya 1, Philip S. Yu1, 1University of Illinois Chicago, 2Salesforce AI Research
1 Introduction
@@
Figure 1: An investigation of the embedding collapse issue of CF model LightGCN he2020lightgcn in two Amazon-review datasets. Increasing embedding size with exponentially growing parameters finally leads to performance degradation.
Recommender systems (RecSys), particularly collaborative filtering (CF), play a pivotal role in many online platforms and web applications  koren2021advances; fan2019graph, aiding users in navigating vast information by offering personalized suggestions. Among the various techniques, collaborative filtering wang2019neural; he2020lightgcn; zhang2024we stands out for its effectiveness in predicting user preferences by analyzing past user-item interactions. A key principle among these models is learning randomly initialized user/item representations with carefully designed loss functions to capture each user’s preference rendle2012bpr. However, such CF approaches heavily rely on the observed user-item interactions to optimize the CF embeddings from random initialization and perform poorly in real-world applications where interaction data is sparse and users/items are cold-start.
More recently, the rapid advancement of large language models (LLMs) such as GPT-4  achiam2023gpt and LLaMa  touvron2023llama, which have exhibited remarkable proficiency in understanding and processing textual information, have piqued the interest of researchers seeking to explore their potential in improving recommender systems beyond traditional approaches  geng2022recommendation; li2023text; zhang2025personaagent. One of the most promising directions in this exploration involves adapting LLMs as recommender systems through prompt engineering and tailored instruction tuning. Such integration of LLMs enables the RecSys to better understand user preferences, and interpret contextual information, thereby surpassing the traditional models in cold-start settings.
@@
Despite the inspiring progress made in LLMs for recommendations, their adoption in online deployments faces significant challenges related to efficiency and scalability. These issues stem from the inherently time-consuming and computationally intensive nature of LLMs. Such problems can be further exaggerated in real-world applications of large-scale users and items due to the enlarged vocabulary tokens and in-context input. Moreover, LLMs fall short of complex CF interaction understanding owing to the next-token prediction pipeline and the constraints of a maximum token limit (e.g., 4096 tokens in LLaMA-2) to incorporate the whole CF interactions. This limitation hampers their ability to capture and effectively model global user-item dependencies as in CF models, leading to poor performance in the full-ranking hou2024large and warm-start settings bao2023tallrec. Therefore, we raise a critical research question:
How can we harness the power of LLMs for RecSys in an effective and scalable manner?
@@
To answer the above question, we propose to utilize LLMs for CF-based RecSys from a new perspective by taking pretrained large language models as a free lunch for the embedding initialization of CF models. One trivial way is to directly utilize the LLMs to generate embeddings with substantial textual information. However, storing the large-scale embeddings from LLMs for each user and item is not efficient nor scalable, especially when the platform scales to millions of users and items. Furthermore, as illustrated in Figure 1, unlike the scaling laws observed in language models kaplan2020scaling and graph models liu2024towards, the performance of the recommendation model LightGCN he2020lightgcn deteriorates even when the numbers of parameters scale up exponentially. This behavior resembles the embedding collapse issue guo2024embedding observed in click-through rate prediction models wang2021dcn. These findings suggest that significantly smaller embedding sizes (e.g., 128) are more suitable for recommendation tasks. In contrast, the top-10 models in the LLM embedding benchmark MTEB muennighoff2023mteb, pretrained for various NLP tasks, use an average large embedding size of 4,506.
Figure 2: An Illustration of LLMInit framework including contextual LLM input, three types of selective item embedding initialization strategies, and the user embedding aggregation operation.
In this paper, we present LLMInit, a practical and scalable initialization framework that bridges large language model (LLM) embeddings with collaborative filtering (CF) models for industrial-scale recommendation systems. LLMInit tackles the deployment challenges of LLMs and enhances the semantic capacity of CF models by initializing CF embedding models based on pretrained LLMs. To this end, we introduce three efficient strategies, including random, uniform, and variance-based index selection, that selectively sample and compress LLM-generated item embeddings to suit the constrained embedding spaces of CF models. On the user side, where contextual signals may be missing or sparse, LLMInit aggregates item-level embeddings using lightweight pooling mechanisms. This design enables recommendation models to inherit rich language semantics from pretrained LLMs while maintaining high efficiency in both training and inference. Extensive evaluations on multiple real-world datasets demonstrate that LLMInit consistently enhances performance across several state-of-the-art CF architectures. Compared to full LLM-based recommendation systems, our framework achieves significant gains in scalability and efficiency, making it well-suited for real-world deployments. Our contributions underscore the value of adapting language model innovations to meet the practical demands of large-scale, industry-grade recommendation systems.
@@
  • Methodologically, we devise three types of selective initialization strategies to inherit the rich pretrained knowledge from LLMs into lightweight CF models for scalable recommendations.
  • Empirically, we implement LLMInit in a plug-and-play manner on various SOTA CF models, achieving significant performance and efficiency gains, especially compared to LLM RecSys.
2 Preliminaries
2.1 Problem Definition
@@
2.2 Related Work
LLMs for Recommendation. Recent advancements in this domain can be categorized into three paradigms wu2024survey: extracting embeddings for sequential recommendation qiu2021u; hou2022towards, generating semantic tokens to capture user preferences li2023text; xi2024towards, and directly adapting LLMs as RecSys through prompting hou2024large; liang-etal-2025-taxonomy or tuning geng2022recommendation; yang2024item. However, they all fail to capture the intricate CF relationship in the user-item bipartite graphs as they either only focus on sequential order patterns qiu2021u; hou2022towards or semantic understandings li2023text; xi2024towards; zhang2023dual. In addition, they face notable challenges in efficiency and scalability due to the computational demands of LLMs, large-scale embeddings, overloaded vocabulary tokens, and limited in-context inputs in real-world scenarios. In contrast, our work LLMInit, proposes a novel paradigm to leverage LLMs in a scalable manner and addresses ineffective interaction modeling by resorting to CF models.
Table 1: Performance comparison on four datasets. Gains are relative to the base model’s performance.
[…]
3 Method
@@
3.1 Selective Item Embedding Initialization
The CF item representation space can be regarded as a subspace of the sophisticated world-knowledge representation space in large language models sheng2024language. Motivated by this, we propose to distill and selectively utilize the embeddings from LLMs to initialize the embeddings for CF recommendation models.
Formally, $\mathbf{v}\in\mathbf{R}^{N}$ represents the item embedding vector generated by the pre-trained LLMs. Let $K$ denotes the desired embedding dimensionality after selection. Then $\mathcal{I}\subseteq\{0,1,\dots,N-1\}$ is the selected indices. The resulting $K$-dimensional embedding is: $\mathbf{v}_{\text{selected}}=\{\mathbf{v}[i]:i\in\mathcal{I}\}.$ Here, we propose three novel initialization strategies to mitigate the inefficiency and embedding collapse issue for directly adopting the LLMs embeddings as follows.
3.1.1 Uniform Selection (LLMInit-Uni)
@@
$\mathcal{I}=\{k\cdot\lfloor N/K\rfloor:k\in\{0,1,\dots,K-1\}\}.$
3.1.2 Random Selection (LLMInit-Rand)
@@
3.1.3 Variance Selection (LLMInit-Var)
@@
3.2 Aggregated User Embedding Initialization
@@
$\mathbf{u}_{i}=\sum_{j\in N_{i}}\frac{1}{{|N_{i}|}}\mathbf{v}_{j},$ (1)
@@
4 Experiments
4.1 Experimental Set Up
We adopt the 5-core setting from zhang2025sgcl; zhang2024mixed and conduct experiments on four real-world Amazon datasets ni2019justifying,  [^1] https://jmcauley.ucsd.edu/data/amazon/links.html, including Beauty, Toys-Games, Tools-Home, and Office-Products, as summarized in Table 4. We use a leave-one-out strategy for data splitting and evaluate all the models using top-K metrics for ranking, including Recall@10 and NDCG@10. For collaborative filtering (CF) baselines, we include LightGCN he2020lightgcn as a representative graph-based model, SGL wu2021self for self-supervised contrastive learning, and SGCL zhang2025sgcl for supervised contrastive learning. For LLM-based baselines, we evaluate MoRec yuan2023go, LLMRank hou2024large, LLMRec wei2024llmrec, and TIGER rajput2023recommender. To ensure fair comparison, we fix the embedding dimension to $K=128$ across all models and use MPNet song2020mpnet as the default embedding generator.
Table 2: Comparisons of computation cost and performance with LLM-based RecSys in full-ranking settings.
@@ 4.1 Experimental Set Up
MoRec yuan2023go 0.0857 0.0397 13M 0.0793 0.0362 14M
LLMRank hou2024large 7B 7B
LLMRec wei2024llmrec 0.0974 0.0472 7B+13M 0.0820 0.0389 7B+14M
TIGER rajput2023recommender 0.1015 0.0481 13M 0.0864 0.0385 13M
LLMInit-Full 0.1029 0.0475 13M 0.0871 0.0387 14M
LLMInit-Var 0.1104 0.0522 2M 0.0941 0.0421 2M
4.2 Performance Comparison
The experimental results clearly demonstrate that all variants of LLMInit substantially improve recommendation performance across diverse datasets compared to traditional CF baselines. Among the proposed strategies, LLMInit-Var consistently achieves the best results, effectively capturing both collaborative filtering signals and semantic information through variance-based index selection. While both LLMInit-Rand and LLMInit-Uni offer notable gains, the variance-based method proves to be the most robust and adaptable across model types and domains. Notably, on the Office-Products dataset (with severe embedding collapse issue as in Figuire 1), LLMInit-Var delivers more than a 20% average gain in NDCG@10 over strong baselines like SGL and SGCL. These consistent improvements highlight the generalizability and practicality of LLMInit, especially in handling diverse real-world scenarios.
4.3 Efficiency and Perforamance Comparisons of LLM-based RecSys